Write the given differential equation in the form L(y) = g(x),

 where L is a linear differential operator with constant coefficients. If possible, factor L. (Use D for the differential operator.)

3y'' − 8y' − 3y = 4


1
Expert's answer
2021-03-30T10:54:44-0400

We know that "y'=Dy, y''=D^2 y," here "D" is differential operator.


"3D^2y-8Dy-3y=4"

"(3D^2-8D-3)y=4"

"(3D+1)(D-3)y=4"